OpenStudy (anonymous):

How does finding the square root [square root pic] of a number compare to finding the cube root [3] of a number? Use the number 64 in your explanation.

OpenStudy (anonymous):

64 = 2^6 = 2^3 2^3 = 2^2 2^2 2^2 64^(1/2} = 2^3 = 8 64^(1/3) = 2^2 = 4

OpenStudy (anonymous):

thanks dude

OpenStudy (anonymous):

Your welcome.